OpenStudy (anonymous):

What is the Y/X intercept of y=3x+3? Please Help!

OpenStudy (anonymous):

x-intercept means y=0 so 3x+3=0 => 3x=-3 => x=-1 y-intercept means x=0 so y=3*0 +3 = 3